OpenStudy (anonymous):

what is the discontinuity of the function f(x) = x^2-3x-28/x+4

OpenStudy (usukidoll):

the denominator will tell you ... all you have to do is find the value of x that would make the denominator 0

OpenStudy (usukidoll):

In this case it's x + 4 = 0 x = -4

OpenStudy (usukidoll):

hold it don't ........ hmmmmmm I didn't see the numerator until now, if we factor that and it becomes x + 4 above then we can cancel

OpenStudy (usukidoll):

and then there will be continuous stuff everywhere...sorry I had a rough night

OpenStudy (campbell_st):

and the discontinuity is a point discontinuity...and not a vertical asymptote as (x + 4) can be removed as a common factor

OpenStudy (usukidoll):

yeah I just factored and found out that the x + 4 can be removed. leaving x - 7 behind

OpenStudy (usukidoll):

and since there's no denominator I don't see how it can be discontinuous. been four years since I touched calc i :P

OpenStudy (campbell_st):

\[f(x) = \frac{(x-7)(x+4)}{(x +4)}\] so really you are left with a straight line f(x) = x - 7 which a point discontinuity

OpenStudy (usukidoll):

maybe point discontinuity when x = 7?

OpenStudy (usukidoll):

blah I should get back to my homework sorry to crash T___T

OpenStudy (campbell_st):

the point discontinuity is at x = -4

OpenStudy (usukidoll):

hmm I was right earlier.. just leave the equation alone and focus on the denominator.

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

None of what you are saying is an option lol

OpenStudy (campbell_st):

well look at the denominator as see what makes it zero x + 4 = 0 this value of x still has relevance ... even though the function can be simplified to a linear function... you could show this be using a table of values... and the original function x: -3 -2 -1 0 1 2 3 y -10 -9 -8 -7 -6 -5 -4 so the table of values shows a linear relationship so what happens when you put x = -4 into the original equation you get y as undefined... but if you put x = -4 into f(x) = x - 7 y = -11 then try x = -5 you get y = -12 so there is a contradiction at x = -4 simplified form it exists... original it doesn't exist hope it helps

OpenStudy (campbell_st):

the points of discontinuity are sometimes called a removable discontinuity here is a link to some notes.. http://classroom.synonym.com/point-discontinuity-algebra-ii-2535.html my best guess on this question... you're expected to factor the numerator remove the common factor then show that the vertical asymptote doesn't exist...

OpenStudy (anonymous):

Yes youve helped thanks :)

OpenStudy (anonymous):

Can you help me with another please :P